OpenStudy (anonymous):

Two numbers have a sum of 24 and a difference of 6. If x represents the larger # and y represents the smaller number, which system and solution represent this information?

OpenStudy (anonymous):

answers could be x + y = 24 y – x = 6 Answer: 15 and 9 x + y = 24 x – y = 6 Answer: 15 and 9 x + y = 6 x – y = 24 Answer: 4 and 2 x + y = 6 x – y = 24 Answer: 15 and 9

OpenStudy (anonymous):

Your answer is: x + y = 24 x – y = 6 Answer: 15 and 9